OpenStudy (anonymous):

express y in terms of x if x/4-3y=7

OpenStudy (anonymous):

x/4=7+3y x/4-7=3y x/12-7/3=y

OpenStudy (anonymous):

set y by itself by first adding over 3y to the other side once done subtract the seven over and then divide by 3 so it should look like\[(\frac{ x }{ 4 }-7)\ (\frac{ 1 }{ 3 })\]

OpenStudy (anonymous):

multiply that through and you get\[\frac{ x }{ 12 }-\frac{ 7 }{ 3 }\]

OpenStudy (anonymous):

x4−3y = 7 ⇒−3y = 7−x4 ⇒−3y = 28−x4 ⇒(−3y)×4 = 28−x ⇒−12y = 28−x ⇒−y = 28−x12 ⇒y = −(28−x12) ⇒y = x−2812 Therefore the required expression on terms of x is y = x−2812.